OpenStudy (anonymous):

You went snow skiing for 6 hours. You skied at an average rate of 30 miles per hour and the ski lift took you back up the hill at a rate of 10 miles per hour. If your overall average rate was 23 miles per hour, how many hours did you actually spend skiing?

OpenStudy (perl):

Let X = time on ski lift. Y = time skiing on slope

OpenStudy (perl):

distance = rate * time X * 10 + Y*30 = 6*23 X + Y = 6